OpenStudy (anonymous):

Can you show me step by step \[\int\limits \frac{ dx }{ 1- x^{4} }\]

OpenStudy (anonymous):

iv'e been having problems with this i think i would have to approach it with u-substitution

OpenStudy (anonymous):

i think i got it if someone could check my answer \[\frac{ 1 }{ 4 }\ln\left| \frac{ 1 + x }{ 1 - x } \right|+\frac{ 1 }{ 2 }\arctan x +c\]